the perimeter of a triangle measures 62cm. the 2nd side is 2 times longer than the fist, and the 3rd side is 12cm larger than the second. What is the length of the longest side.
Answer by pwac(253)

Ok
Let first side = x
so second side is 2x
and third side is 2x+12
so now we have x+2x+2x+12=62 sum the x-terms
5x+12=62 now subtract 12 from both sides of the equation
5x+12-12=62-12
5x=50 divide both sides by 5 to get
x=10 this is length of first side
second side is 2x which =20
third side is 2x+12 which =32
so 10+20+32 = 62cm
Pete